PRM oplev beam was not hitting on the QPD since Jun 1, so I centered it. I reverted the oplev servo gains and now oplev servo looks fine.
C1:SUS-PRM_OLPIT_GAIN = 1.0
C1:SUS-PRM_OLYAW_GAIN = -0.7
There's SIDE to UL/UR/LL/LR coil element in PRM TO_COIL matrix. They were 0 until Mar 31, 2012, but someone changed them to -0.160. I couldn't find elog about it.
Same thing happened to BS on Mar 13, 2012 (see elog #6409), so I think somebody did the same thing to PRM. |